OpenStudy (anonymous):

Solve the following quadratic equation by completing the square 7x^2+5x+72=3x^2-19x

OpenStudy (calculusxy):

http://youtu.be/xGOQYTo9AKY

OpenStudy (texaschic101):

get all the single numbers on the right side... 7x^2 + 5x + 72 = 3x^2 - 19x 7x^2 - 3x^2 + 5x + 19x = -72 combine like terms 4x^2 + 14x = -72 Still with me ?

OpenStudy (anonymous):

ok i got you

OpenStudy (anonymous):

but 24x instead of 14x right

OpenStudy (texaschic101):

Now take half of your x, which is 7, and square it. (7 x 7 = 49) 4x^2 + 14x + 49 = -72 + 49 combine like terms 4x^2 + 14x + 49 = - 23 still with me ?

OpenStudy (texaschic101):

oops...wait...mistake ..hold on

OpenStudy (texaschic101):

4x^2 + 24x = -72 4x^2 + 24x + 144 = -72 + 144 combine like terms 4x^2 + 24x + 144 = 72 understand so far ? I got the 144 by taking half of coefficient of x (24) which is 12 and then I squared 12 and added to both sides.

OpenStudy (anonymous):

yes i understand

OpenStudy (anonymous):

Thats it?

OpenStudy (texaschic101):

no...hold on, this is where I have problems....give me a sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

should divide by four to get the x^2 by itself?

OpenStudy (texaschic101):

I should have factored.... 4x^2 + 24x + 72 = 0 4(x^2 + 6x + 18) = 0 now we have x^2 + 6x +18 = 0 x^2 + 6x = -18 x^2 + 6x + 9 = -18 + 9 x^2 + 6x + 9 = -9 (x + 3)^2 = -9 I am sorry.....you will need to get better help....I am not good at these....tag someone that can really help you. Again, I am sorry

OpenStudy (texaschic101):

@satellite73 ...can you help here....I am so not good at these

OpenStudy (texaschic101):

@hartnn ...can you help...I am not good at this stuff

hartnn (hartnn):

ok, are we clear till : get all the single numbers on the right side... 7x^2 + 5x + 72 = 3x^2 - 19x 7x^2 - 3x^2 + 5x + 19x = -72 so, 4x^2 +24x =-72 ?

hartnn (hartnn):

then divide by 4 throughout to get x^2+6x = -12 ?

OpenStudy (anonymous):

i got 18 instead of 12

hartnn (hartnn):

oops, sorry x^2+6x =-18\ yes

hartnn (hartnn):

now step 1 : divide the co-efficient of x by 2 whats the co-efficient of x here ?

OpenStudy (anonymous):

-1/6?

hartnn (hartnn):

for ax^2+bx=c the co-efficient of x is just "b" so, here, the co-efficient of x is just 6, got this ?

OpenStudy (anonymous):

ok

hartnn (hartnn):

divide it by 2 , what u get ?

OpenStudy (anonymous):

x^2+2x=-9 or does the -18 stay the same

hartnn (hartnn):

how did u get the 2x ?

OpenStudy (anonymous):

i mean 3x

hartnn (hartnn):

not actually.... i just asked to divide 6 by 2 no modification in original equation yet. it still stays as x^2+6x = -18

hartnn (hartnn):

so, since 6/2 = 3, is the step 1 output, lets got to step 2: Square it. so whats 3^2 =...?

OpenStudy (anonymous):

9

hartnn (hartnn):

good, so lets take the original equation now step 3: add that to both sides of the equation. so, what do u get when you add 9 on both sides of x^2+6x =-18 ?

hartnn (hartnn):

u still there ?